Two spheres, P and Q with charges of +4 microcoulomb and -6 microcoulomb are brought into contact and separated again.
What is the charge on each sphere after contact?
1
Expert's answer
2020-05-28T13:07:27-0400

The potential of the first sphere before after contact is:


"\varphi_1' = \dfrac{1}{4\pi\varepsilon_0}\dfrac{q_1'}{R_1}"

where "q_1'" is the charge on the first sphere after contact and "R_1" is the radius of the first sphere.

The potential of the second sphere after contact is:


"\varphi_2' = \dfrac{1}{4\pi\varepsilon_0}\dfrac{q_2'}{R_2}"

where "q_2'" is the charge on the second sphere after contact and "R_2" is the radius of the second sphere.

As far as the charge conserves, the total anoum of charge will not change:

"q_1' + q_2' = 4\mu C + (-6\mu C) = -2\mu C"


After contact the potentials of the spheres will be equal to each other:


"\varphi_1' = \varphi_2' \Rightarrow \dfrac{q_1'}{R_1} = \dfrac{q_2'}{R_2}"

Express "q_1'" and substitute it to the previous equation:

"q_1' = q_2'\dfrac{R_1}{R_2}\\
q_2'\dfrac{R_1}{R_2} + q_2' = -2\mu C\\
q_2' = -2 \dfrac{R_2}{R_1 + R_2} \mu C"

The charge of the first sphere will be:


"q_1' = -2 \dfrac{R_1}{R_1 + R_2} \mu C"